# SeagrassFinder: Deep Learning for Eelgrass Detection and Coverage Estimation in the Wild

### v1

@@ II-B Data Annotation and Platform
The interface of SeagrassFinder AP is designed to be straight-forward with minimal amount of text to keep the barrier of entry for a new user as low as possible (see Fig. 3(a)). We also keep the explanation extremely brief and simplify the task instructions by adding visual guidance with concrete image examples with eelgrass present and absent. Alongside the two label buttons we add a button to skip the current sampled image as well as a button to undo the previous annotation. A user can also submit a comment if an image is considered invalid. For each annotation task we uniformly sample an image from the data set and display it to the annotator to maximise objectivity and to avoid showing consecutive frames which could potentially add labeling bias. To motivate users to annotate as many images as possible we add a leaderboard to the platform, to spark competitiveness between users which proved highly effective.
@@
Figure 3: (a) A screenshot of the SeagrassFinder Annotation Platform accessed through a browser window. The labeling interface is simple and intuitive. (b) A screenshot of the start page of the SeagrassFinder AP showing the visual guidance for the labeling instructions.
II-C Annotation Analysis
We conduct an analysis of the annotations collected via SeagrassFinder AP since their correctness has a direct influence on the performance of our DNN model predictions. In total we have 19 different users creating more than 13000 annotations on more than 8900 different images. We calculate metrics to analyze the inter-annotator agreement level such as annotation agreement percentage, Cohen’s Kappa for reliability of annotators as well as mistake rate, i.e. rates of “mistake” and “ambiguous” annotations.
Generally, all annotators exhibit a high degree of annotation agreement, with annotators agreeing with one another on average 76% of the time. Very few annotators (3 out of 19) exhibit a lower than 70% agreement. We also investigate the impact of non-expert annotators annotating images and found that no bias is present between images annotated by non-expert annotators. There is also no connection between annotators with a lower agreement level, and their domain experience. This confirms that the classification of UW images for eelgrass presence is an annotation task humans are capable of, independent of annotators being experts or non-experts. This is still true for more challenging images with high levels of blur, color-distortion or low contrast.
We report how many times an image is labeled in Fig. 4(a). Additionally we calculate the probability of agreement for each image normalized relative to the number of annotations (see Fig. 4(b)). For images that are annotated twice, the annotators achieve an overall 93% agreement. Cohen’s Kappa scores are calculated between every annotator’s subset of common annotated images and reported in the format of a matrix (see appendix Fig. 11). Using magnitude guidelines defined by Landis et al. (Landis, 1977) we find that on average there is a level of agreement is of substantial to almost perfect magnitude. In occasions where less than or moderate agreement is seen, we find that this is caused by a too small sample size of commonly annotated images.
To estimate the mistake rate of annotators we consider two types of disagreements in annotations which we verify manually; a “mistake annotation” where a user accidentally clicks the wrong label button caused by an obvious lapse in judgment by visual inspection, and an “ambiguous annotation” where eelgrass is only partially visible or eelgrass presence was uncertain. By labeling ambiguous annotations we estimate to which degree annotators make mistakes, and if the annotation task was not defined well enough. Based on the ratio of ambiguous to mistake annotations we dismiss the presumption that the annotation task was too vague. However it does highlight that the task of annotating can be at times very difficult to complete. Among the annotators, mistake annotations occur at a fairly low rate, on average at around 8% (see Fig. 12.) There seems to be no relation between the level of disagreement percentage of a user and their mistake percentage.
II-D Final Dataset Compilation
@@
A detailed overview of our data set strategy for the DNN training and testing pipeline can be seen in Tbl. II. First we manually remove all images taken above sea-level manually to mitigate inclusion of non-relevant features (this can be done quite quickly by trimming the video to once it is underwater). Additionally we filter out all images with disagreeing annotations. The final dataset consists of 8324 images, 52% of which are labeled as “Eelgrass Present”. To avoid data leakage between train and test set we do not mix images between different transects. Throughout the project, we hold out transect LYT-9 as the final test dataset. In all experiments we split the train data into validation sets of 20% using random split procedure.
@@
Transect Name #Images #Images Annotated Eelgrass Present Init. Train Dataset Init. Test Dataset Cross Val. Dataset Final Train Dataset Final Test Dataset
@@
TABLE II: A summary of our data strategy: each transect results in a video from which we extract and annotate a subset of images. We show the percentage of images with eelgrass labels. We list the datasets used for initial hyperparameter tuning and cross validation (cross val.). We excluded LYT-5 and LYT-14 from cross validation because of their high class imbalance. LYT-9 was chosen as the final test set and therefore excluded from all other training/testing procedures. All remaining images were used in the final model training.
II-E Deep Learning Methodology
In this work, we adopt transfer learning to train a number of DNN models on the specific task of classifying images for two classes; “Eelgrass Present” and “Eelgrass Absent”. We analyse four different model architectures;
@@ II-E Deep Learning Methodology
All models are available online with pre-trained weights in PyTorch (Paszke et al., 2019), and are pre-trained on the ImageNet (Russakovsky et al., 2015) dataset. We fuse each model with two additional fully-connected layers for performing classification. During the training process, we only fine-tune the weights of the additional layers while the weights of each pre-trained model are frozen. A visualization of the transfer learning mechanism can be seen in Fig. 5. We evaluate the performance of our chosen DNN models according to three metrics:
@@
Throughout this document, where higher values mean better performance, metrics are marked with ($\uparrow$) and where lower values mean better performance metrics are marked with ($\downarrow$).
Figure 5: Transfer learning: (1) we use DNN models pre-trained on a generic data set, here ImageNet (Russakovsky et al., 2015). The trained layers of each DNN model are frozen and infused into our custom training process (2). We fine-tune a small number of additional layers on our eelgrass data set to perform the target task of eelgrass detection (3).
We choose four DNN models in this work to scrutinize their ability on our custom classification task. Generally we want to test and analyze a range of different architectures, but also compare models previously mentioned in the literature in this context. Our end goal is to identify the DNN architecture that can perform our target task of eelgrass detection with the highest accuracy. The ResNet models are well-known in their performance capabilities in image classification (Tan et al., 2018). InceptionNetV3 is mentioned previously in the literature (Reus et al., 2018) and its common use in image classification tasks. DenseNet architecture models show high accuracy in other image classification contexts, such as medical imaging (Kim et al., 2022), and efficiency in training time compared to ResNet architectures. Finally we choose the Vision Transformer model (configuration ViT_l_32) because they are considered state-of-the-art in image classification. We believe we are the first to apply ViTs in the context of eelgrass classification.
One fundamental challenge in assessing underwater image data is the light attenuation in water. To mitigate these factors and to increase visual quality of the images, we employ a DNN-based underwater (UW) image enhancement tool called DeepWave-Net (Sharma et al., 2023). We are curious whether this could have a positive effect on prediction accuracy of our DNN models because the images of ImageNet are predominantly taken on land, and thus do not exhibit above-mentioned shortcomings. We conduct all model training and testing procedures with and without the application of UW-enhancement (see Figure 6). The UW-enhanced images show a wider range of colors, a better contrast between the green and red channels on the RGB color range. This enables easier differentiation between the green/brown filamentous algae from the eelgrass itself, as well as the sugar kelp within the image. Furthermore, when applying the model on images at a greater water depth, we notice that the effect is even more significant.
@@
To evaluate each model architecture, we run multiple machine learning experiments. To gauge baseline performance of each model architecture, we complete an initial experiment. Here we choose to use LYT-20 as a test dataset, due to its frequent changing underwater landscape, caused by it’s zig-zag constellation (see Fig. 2(d)). For training, we use all other transects, except LYT-9. LYT-9 is excluded since it is used as the final test dataset, and must be excluded for all experiments. For these experiments we create the validation dataset by using a random split of a 20% subset of images from the train dataset. The DNN classification layer configuration for the initial run (last layer and second last layer size) are configured as seen in Tbl. III. For these runs we choose to configure the second last layer with more neurons than the last layer to guide the learning process towards a natural dimensionality reduction.
@@
After completion of the initial evaluation experiments, we choose a subset of models which performed best, and complete a cross-validation test. To overcome the challenge of imbalances in class distributions between different transects we devise an alternative approach to cross validation compared to traditional k-fold cross validation. We call this leave-one-transect-out cross validation. This approach only uses data from transects LYT-10, LYT-12 and LYT-20 due to their similar class distributions. On each cross validation fold we choose one of the three transects as the validation set, and the model trained on all other transects. We repeat this process of selecting one transect for validation three times, as such allowing for training and testing across all three transects. Upon completion we average the validation scores taken from all three experiments. This gives a good estimation of model performance, which is not directly overfit to one specific train and test dataset. The hyperparameters chosen for these experiments are the same as the hyperparameters used in initial experiments. Finally, based on the “leave-one-transect-out” validation test results, we select the two highest performing models and complete the final model experiment. This involves using all transects except LYT-9 as the train/validation dataset and test on the LYT-9 transect.
Model Last Layer Second To Last Layer #Trainable Weights #Frozen Weights Learning Rate
ResNet 512 512 131K 11.4M 1.7e-05
InceptionNetV3 256 512 1.2M 25.1M 1.7e-05
DenseNet 256 512 623K 18.1M 1.7e-05
ViT 256 512 513K 306M 1.7e-05
TABLE III: Model Hyperparameters: For each chosen model architecture we report the size of the two additional layers (last and second to last layer) to be fine-tuned by transfer learning, the number of trainable weights for fine-tuning, the number of frozen weights and the learning rate.
II-F Eelgrass Coverage Estimation and Postprocessing
In a traditional eelgrass transect methodology eelgrass coverage is defined by having marine biologists annotate how much of the observed area is covered in eelgrass as a percentage. In Denmark eelgrass coverage is defined as “the total substrate-specific coverage (determined) by projecting the outline of the foliage vertically onto the surface of the soft bottom and assessing the foliage’s percentage coverage of the bottom.” - translated from Danish (Bruhn et al., 2013). We find that this definition is unpractical to comply with when considering EIA video transects, where images are taken at an angle towards the seafloor. Alternatives such as visual coverage guides are highly subjective; the difference for example between 30% and 55% is extremely hard to determine (Short et al., 2015).
In this work we propose steps towards a novel, data-driven method for estimating eelgrass coverage. We are interested in developing a coverage estimation method that is based on the frequencies of outputs of the sample-based binary presence predictions of an image stream containing varying levels of eelgrass abundance. Since we consider spatially overlapping frames, we estimate the area with eelgrass present along the transect by calculating the temporal mean of the image stream with 0% representing eelgrass absent and 100% representing eelgrass present. The general expression of the temporal mean (TM) in a sequence of predictions $p$ in a range $r$ at instance $i$ is given by:
$\textrm{TM}_{r,i}=\frac{p_{i}+p_{i+1}+p_{i+2}+\ldots+p_{i+r-1}}{r}=\frac{1}{r}\sum^{n=r-1}_{n=0}p_{i+n}.$ (1)
Figure 7: Limitations: even though the level of eelgrass abundance is different in the top scenes compared to the bottom scenes, eelgrass coverage estimation based on a temporal mean of ten samples gives the same result in both examples.
There are some limitations to our coverage estimation method as depicted in Fig. 7. First, consecutive images that have consistently little eelgrass present give the same coverage estimation as consecutive images with abundant eelgrass (example 1). Therefore, there is no distinction in the abundance as such, since the frequency of images with the binary prediction “eelgrass present” is the same. Second, it is not possible to determine the location of small and dense eelgrass patches as they give the same coverage estimation as larger sparse patches. The temporal mean obliterates this information (example 2). Third, the method is sensitive to the temporal mean window size. Even though these limitation exist we argue that they are outweighed by the advantages of our method mentioned above.
To extract the overlaid positional meta-data from the transect videos we use the Tesseract Engine (Smith, 2007) for optical character recognition. We fuse the resulting GPS data with our eelgrass annotations and DNN-based predictions. We choose transects LYT-9, LYT-10, LYT-11, LYT-12, LYT-13, LYT-19 and LYT-20. For each transect, we use our ViT model to get a prediction for every 10th frame. Since all the videos are filmed at 30 frames per second, this gives three predictions per second. In total this creates 16796 predictions of eelgrass presence, related to a geographical position along each of the transects. To further give intuition on the benefits of our novel estimation coverage approach we conduct a visual comparison between our approach and human annotations on a transect video from an EIA in a different location in Denmark (around 150km away), since we have annotated coverage estimations available for this specific video transect. We cut a 15 minute section (around 900m to 1350m) from this transect and use our ViT predictions to estimate coverage.
III Results
III-A Initial Model Evaluation

### v2

@@ 2.2 Data Annotation and Platform
The interface of SeagrassFinder AP is designed to be straight-forward with a minimal amount of text to keep the barrier of entry for a new user as low as possible (see Fig. 4). We also keep the explanation extremely brief and simplify the task instructions by adding visual guidance with concrete image examples, both with eelgrass present and absent. Alongside the two label buttons we add a button to skip the current sampled image as well as a button to undo the previous annotation. A user can also submit a comment if an image is considered invalid. For each annotation task we uniformly sample an image from the data set and display it to the annotator to maximize objectivity and to avoid showing consecutive frames which could potentially add labeling bias. To motivate users to annotate as many images as possible we add a leaderboard to the platform, to spark competitiveness between users which proved highly effective.
@@
Figure 4: A screenshot of the SeagrassFinder Annotation Platform accessed through a browser window. The labeling interface is simple and intuitive.
2.3 Dataset Compilation and Training Strategy
@@
A detailed overview of our data set strategy for the DNN training and testing pipeline can be seen in Tbl. 2. First we manually remove all images taken above sea-level manually to mitigate inclusion of non-relevant features. Additionally we filter out all images with disagreeing annotations. The final dataset consists of 8324 images, 52% of which are labeled as “Eelgrass Present”. To avoid data leakage between train and test sets we do not mix images between different transects. Throughout the project, we hold out transect LYT-9 as the final test dataset. In all experiments we split the training data into validation sets of 20% using random splits.
Table 2: SeagrassFinder data strategy: each transect results in a video from which we extract and annotate a subset of images. We list the percentage of images with eelgrass labels, the datasets used for initial hyperparameter tuning and cross validation (cross val.). We excluded LYT-5 and LYT-14 from cross validation because of their high class imbalance. LYT-9 was chosen as the final test set and therefore excluded from all other training/testing procedures. All remaining images were used in the final model training.
@@
Transect Name #Images #Images Annotated Eelgrass Present Initial Train Set Initial Test Set Cross Val. Set Final Train Set Final Test Set
@@
2.4 Data Preprocessing
One fundamental challenge in assessing underwater image data is the light attenuation in water with increasing water depth. To mitigate these factors and to increase visual quality of the images, we employ a DNN-based underwater (UW) image enhancement tool called DeepWave-Net (Sharma et al., 2023). We are curious whether this could have a positive effect on prediction accuracy of our DNN models. We conduct all model training and testing procedures with and without the application of UW-enhancement (see Figure 5). The UW-enhanced images show a wider range of colors, a better contrast between the green and red channels on the RGB color range. This enables easier differentiation between the green/brown filamentous algae from the eelgrass itself, as well as the sugar kelp within the image. We notice that the effect is even more noticeable, when applying the model on images at a greater water depth.
2.5 Deep Learning for Eelgrass Detection
Figure 6: Transfer learning: (1) we use DNNs pre-trained on a generic data set (blue), here ImageNet (Russakovsky et al., 2015). (2) The trained layers of each DNN model are frozen and infused into our custom training process. (3) We fine-tune a small number of additional layers on our eelgrass data set (green) to perform the target task of eelgrass detection.
In this work, we adopt transfer learning to train a number of DNN models on the specific task of classifying images for two classes; “Eelgrass Present” and “Eelgrass Absent”. In transfer learning, a DNN previously trained on a large generic data set, such as ImageNet (Russakovsky et al., 2015) for the task of classification, is reused for the same task but on a different, typically more specialized and smaller data set. For example, ImageNet-1k contains 1000 classes of animals and common objects, e.g. cats, dogs and cars. We are interested in the process of fine-tuning; instead of starting the DNN training process from scratch, we freeze the DNN layers trained to classify ImageNet, and add additional layers to the DNN architecture. By only training these additional layers on our specialized task of distinguishing between eelgrass absence and presence, we can reuse the DNN’s knowledge and understanding of the generic task. Not only does this save computational resources, it also drastically improves learning efficiency and typically classification accuracy. Transfer learning is a popular approach in deep learning, to overcome data sparsity for specialized machine vision tasks. A visualization of the transfer learning mechanism can be seen in Fig. 6.
One class of DNNs are convolutional neural networks. These contain convolutional layers that can learn features, like edges, corners and textures, by convolving multiple learnable filters across an input image. The output of these convolutional layers, a set of features maps, are the local responsiveness of an image area to each filter, while keeping the spatial relationships consistent. Convolutional neural networks are efficient at learning representative features, contributing to their effectiveness in image classification and object detection. However, they have inductive bias; they inherently assume that pixels in proximity have local dependencies that can be exploited during the learning process. This may not always be the case where a larger spatial integration is required for better contextual and spatial understanding. The current state-of-the-art DNN is the Vision Transformer model (ViT) (Dosovitskiy, 2020). These can overcome inductive biases by assigning a positional encoding to image patches and by incorporating a global attention mechanism, inspired by the latest advancements in natural language processing.
We choose four DNN models in this work to analyze their ability to classify seagrass presence in images. We aim to test and analyze a range of state-of-the-art DNN architectures and model sizes. ResNet models are widely recognized for their strong performance capabilities in image classification tasks (Tan et al., 2018). InceptionNetV3 is a well-established architecture and also referenced in prior work in seagrass analysis (Reus et al., 2018) and its popularity in image classification tasks. DenseNet architecture models show high accuracy in other image classification domains, such as medical imaging (Kim et al., 2022). Finally we choose Vision Transformers (configuration ViT-l-32) which are widely considered state-of-the-art in image classification. Based on our review, we are the first to apply ViTs in the context of eelgrass classification. Since our primary focus in this work is benchmarking model performance rather than proposing new architectures, we refer interested readers to the following works for in-depth introductions to convolutional neural networks and Vision Transformer architectures, as well as comparative evaluations in related domains such as medical imaging (Zhao et al., 2024; Taher et al., 2025; Takahashi et al., 2024). The selected architectures are summarized below;
@@ 2.5 Deep Learning for Eelgrass Detection
All models are available online with pre-trained weights in PyTorch (Paszke et al., 2019), and are pre-trained on the ImageNet (Russakovsky et al., 2015) dataset. We fuse each model with two additional fully-connected layers for performing classification. During the training process, we only fine-tune the weights of the additional layers, while the weights of each pre-trained model are frozen.
Table 3: Model Hyperparameters: For each chosen model architecture we report the size of the two transfer learning (TF) layers for fine-tuning, the number of trained weights for fine-tuning, the number of frozen weights and the learning rate.
Model 1st TF Layer 2nd TF Layer #Trained Weights #Frozen Weights
ResNet 512 512 131K 11.4M
InceptionNetV3 512 256 1.2M 25.1M
DenseNet 512 256 623K 18.1M
ViT 512 256 513K 306M
2.6 Eelgrass Coverage Estimation
In a traditional eelgrass transect methodology eelgrass coverage is defined by having marine biologists annotate how much of the observed area is covered in eelgrass as a percentage. In Denmark eelgrass coverage is defined as “the total substrate-specific coverage (determined) by projecting the outline of the foliage vertically onto the surface of the soft bottom and assessing the foliage’s percentage coverage of the bottom.” - translated from Danish (Bruhn et al., 2013). We find that this definition is unpractical to comply with when considering EIA video transects, where images are taken at an angle towards the seafloor. Alternatives such as visual coverage guides are highly subjective; the difference for example between 30% and 55% is visually hard to determine (Short et al., 2015).
In this work we propose steps towards a novel, data-driven method for estimating eelgrass coverage. We are interested in developing a coverage estimation method that is based on the frequencies of outputs of the sample-based binary presence predictions of an image stream containing varying levels of eelgrass abundance. We bear in mind that our image data has certain characteristics; the images are taken at an angle towards the seafloor and thus the fields of view of two consecutive frames have spatial overlap. We speculate that we can therefore estimate the eelgrass coverage along the transect by calculating the temporal mean of the image stream with 0% representing eelgrass absent and 100% representing eelgrass present. The general expression of the temporal mean (TM) in a sequence of predictions $p$ in a range $r$ at instance $i$ is given by:
$\centering\textrm{TM}_{r}(i)=\frac{p_{i}+p_{i+1}+p_{i+2}+\ldots+p_{i+r-1}}{r}=\frac{1}{r}\sum^{n=r-1}_{n=0}p_{i+n}.\@add@centering$
(a) Example 1
(b) Example 2
Figure 7: Limitations: even though the level of eelgrass abundance is different in the top scenes compared to the bottom scenes, eelgrass coverage estimation based on a temporal mean (TM) of ten samples gives the same result in both examples.
There are some limitations to our coverage estimation method as depicted in Fig. 7. First, consecutive images that have consistently little eelgrass present can give the same coverage estimation as consecutive images with abundant eelgrass (Fig. 7(a)). Second, it may not be possible to determine the exact location of small and dense eelgrass patches as they give the same coverage estimation as larger sparse patches. The temporal mean obliterates this information (Fig. 7(b)). Third, the method is sensitive to the temporal mean window size. Even though these limitations exist we argue that they are outweighed by the advantages of our method.
To extract the overlaid positional meta-data from the transect videos we use the Tesseract Engine (Smith, 2007) for optical character recognition. We fuse the resulting GPS data with our eelgrass annotations and DNN-based predictions. We choose transects LYT-9, LYT-10, LYT-11, LYT-12, LYT-13, LYT-19 and LYT-20. For each transect, we use our ViT model to get a prediction for every 10th frame. Since all the videos are filmed at 30 frames per second, this gives three predictions per second. In total this creates 16796 predictions of eelgrass presence, related to a geographical position along each of the transects.
3 Experimental Design
The overall objective for our experimental design is to identify the best performing and most robust DNN model out of our selection for eelgrass presence detection, such that we can subsequently employ its predictions for postprocessing and eelgrass coverage estimation. Since the quality of our training data is crucial we also conduct an annotation quality assessment. Our data strategy is shown in Tbl. 2. We evaluate the performance of our chosen DNN models according to three metrics;
@@ 3 Experimental Design
In this document, where higher values mean better performance, metrics are marked with ($\uparrow$) and where lower values mean better performance metrics are marked with ($\downarrow$).
3.1 Annotation Quality Assessment
We conduct an analysis of the annotations collected via SeagrassFinder AP since their correctness has a direct influence on the performance of our DNN model predictions. We calculate metrics to analyze the inter-annotator agreement level such as annotation agreement percentage, Cohen’s Kappa for reliability of annotators as well as mistake rate. To estimate the mistake rate of annotators we consider two types of disagreements in annotations which we verify manually; a “mistake annotation” where a user accidentally clicks the wrong label button caused by an obvious lapse in judgment by visual inspection, and an “ambiguous annotation” where eelgrass is only partially visible or eelgrass presence was uncertain. By labeling ambiguous annotations we estimate the degree to which annotators make mistakes, and if the annotation task was not defined well enough. We additionally report the number of times an image is labeled and calculate the probability of agreement for each image normalized relative to the number of annotations.
3.2 Initial Model Evaluation
@@
To gauge the baseline performance of each model architecture, we run an initial experiment where we choose LYT-20 as a test dataset, due to its frequent changing underwater landscape, caused by it’s zig-zag constellation (see Fig. 3(d)). For training, we use all other transects, except LYT-9. For these experiments we create the validation dataset by using a random split of a 20% subset of images from the train dataset. The DNN classification layer configuration for the initial run (last layer and second last layer size) are configured as seen in Tbl. 3. For these runs we choose to configure the second last layer with more neurons than the last layer to guide the learning process towards a natural dimensionality reduction. All models are trained at a learning rate of $4.7\times 10^{-5}$, 33 epochs, batch size 22 and an early stopping mechanism with patience of 5 epochs.
3.3 Leave-One-Transect-Out Validation
@@
We choose a subset of models which performed best, and complete a cross-validation test. To overcome the challenge of imbalances in class distributions between different transects we devise an alternative approach to cross validation compared to traditional k-fold cross validation. We call this leave-one-transect-out cross validation. This approach only uses data from transects LYT-10, LYT-12 and LYT-20 due to their similar class distributions. On each cross validation fold we choose one of the three transects as the validation set, and the model trained on all other transects. We repeat this process of selecting one transect for validation three times, as such allowing for training and testing across all three transects. Upon completion we average the validation scores taken from all three experiments. This gives a good estimation of model performance, which does not directly overfit to one specific train and test dataset. The hyperparameters chosen for these experiments are the same as the hyperparameters used in initial experiments.
3.4 Final Model Evaluation
Finally, based on the results of our “leave-one-transect-out” validation, we select the two highest performing models and complete the final model experiment. This involves using all transects except LYT-9 for training and validation and subsequently test on the LYT-9 transect.
3.5 Eelgrass Coverage Estimation
Since there are no expert labels available of eelgrass coverage estimation for the Lynetteholm EIA, we are limited to giving an intuition of the benefits of our novel estimation coverage. First, we visualize eelgrass presence predictions geographically alongside our eelgrass coverage estimations as a heatmap. To provide further evidence, we show a visual comparison between our approach and human annotations on a transect video from a different EIA in another location in Denmark (around 150km away). Here we have expert annotated coverage estimations. We cut a 15 minute section from this transect and show a visual comparison of our coverage predictions and expert labels in a time series.
4 Results
In this section we present our results in the order of our data processing pipeline “SeagrassFinder”. First we present our annotation quality analysis. Then we evaluate our chosen DNNs according to the task of binary classification of eelgrass presence and absence. Based on these results we choose the best-performing DNN and use its predictions for eelgrass coverage estimation.
4.1 Annotation Quality Assessment
In total we have 19 different users creating more than 13000 annotations on more than 8900 different images. Generally, all annotators exhibit a high degree of annotation agreement, with annotators agreeing with one another on average 76% of the time. Very few annotators (3 out of 19) exhibit a lower than 70% agreement. We also investigate the impact of non-expert annotators annotating images and found that no bias is present between images annotated by non-expert annotators. There is also no connection between annotators with a lower agreement level, and their domain experience. This confirms that the classification of UW images for eelgrass presence is an annotation task humans are capable of, independent of annotators being experts or non-experts. This is still true for more challenging images with high levels of blur, color-distortion or low contrast.
Figure 8(a) shows the number of times an image is labeled. The probability of agreement for each image normalized relative to the number of annotations can be seen in Fig. 8(b). For images that are annotated twice, the annotators achieve an overall 93% agreement. Cohen’s Kappa scores are calculated between every annotator’s subset of common annotated images and reported in the format of a matrix (see appendix Fig. 11). Using magnitude guidelines defined by Landis et al. (Landis, 1977) we find that on average there is a level of agreement that is of substantial to almost perfect magnitude. On occasions where less than or moderate agreement is seen, we find that this is caused by too small a sample size of commonly annotated images.
Based on the ratio of “ambiguous” to “mistake” annotations we dismiss the presumption that the annotation task was too vague. However it does highlight that the task of annotating can be at times very difficult to complete. Among the annotators, mistake annotations occur at a fairly low rate, on average at around 8% (see Fig. 12.) There seems to be no relation between a user’s percentage level of disagreement and their mistake percentage.
4.2 Initial Model Evaluation